OpenStudy (anonymous):

How to factor (2x^2-9x-5). Preferably not the answer, simply an explanation.

OpenStudy (anonymous):

2x^2 + 9x - 5 = 0 (2x -1)(x+5) = 0 2x - 1 = 0 x+5 = 0 x = 1/2 x=-5

OpenStudy (anonymous):

How does multiplying a by c get rid of a Ventricate?

OpenStudy (anonymous):

sorry I meant multiply 2 to 5, so you would have x^2- 9x-10, but when you factor you add the coefficient 2 that you multiplied at first resulting in (2x-10) (2x-1) and solve from there. Sorry for the super late reply!